if you have a mixture of 80g ethanol and 20g methanol. calculate mole fractions of each
Answer #1

Molar mass of C2H5OH,

MM = 2*MM(C) + 6*MM(H) + 1*MM(O)

= 2*12.01 + 6*1.008 + 1*16.0

= 46.068 g/mol

Molar mass of CH3OH,

MM = 1*MM(C) + 4*MM(H) + 1*MM(O)

= 1*12.01 + 4*1.008 + 1*16.0

= 32.042 g/mol

n(C2H5OH) = mass/molar mass

= 80.0/46.068

= 1.7366

n(CH3OH) = mass/molar mass

= 20.0/32.042

= 0.6242

n(C2H5OH),n1 = 1.7366 mol

n(CH3OH),n2 = 0.6242 mol

Total number of mol = n1+n2

= 1.7366 + 0.6242

= 2.3607 mol

Mole fraction of each components are

X(C2H5OH) = n1/total mol

= 1.7366/2.3607

= 0.736

X(CH3OH) = n2/total mol

= 0.6242/2.3607

= 0.264

Answer:

Mole fraction of ethanol = 0.736

Mole fraction of methanol = 0.264